A merry-go-round of radius R = 3.00 m has a moment of inertia I = 250 kg m2 and is rotating at 10.0 rev/min about a vertical axis with friction (coefficient of 0.3). Facing the axle, a 30.0 kg child hops onto the merry go-round and manages to sit down on the edge.

(a) What is the new angular speed of the merry-go-round?
(b) What would the angular speed the merry-go-round be if the child would sit at a distance of 1.0 m from the axle?
Answer #1

using conservation of angular momentum

I w1 = ( I + Mr^2) w2

250 * 10 = ( 250 + 30* 3^2) w2

w2 = 4.808 rev/min

==========

b)

w2 = 25* 10 / ( 250 + 30* 1)

w2 = 8.93 rev/min

========

Comment in case any doubt, will reply for sure. Goodluck